# Does rare, noise-induced, bypass transition in plane Couette flow bypass instantons ?

### v1

@@ 3.1.3 Bisections
Following the procedure presented in section 2.2.4, we perform relaxations and bisections using reactive trajectories computed by means of AMS. We will examine the type of unstable attractor (unstable fixed point, unstable limit cycle, unstable travelling wave) which is reached and describe its main features. The result of these bisections will be compared to reactive trajectories and the last state at the last stage examined in the former subsection. We will check whether the system displays a standard instanton phenomenology, which would be characterised by greater and greater correspondence between $\mathbf{u}_{\rm last}$ and the result of bisections: this would mean the trajectories would cross the separatrix very near the saddle. The time series of these relaxations and the result of following bisections are added to the plots:
@@
This rescaling matches with a $x$ position. It is incremented by $L_{x}$ each time $\arg(\hat{u}_{z})$ goes from $2\pi$ to 00 and decremented by $-L_{x}$ each time $\arg(\hat{u}_{z})$ goes from 00 to $2\pi$. The succession of plateau separated by a rapid jump every period of the kinetic energy quantifies the fact that we periodically see the pattern in $\omega_{x}$ disappear and reappear at a distance $L_{x}/2$. This edge state in the form of an unstable limit cycle has not been specifically studied in the literature to our knowledge, possibly because of the smallness of the domain. However, a similar spatial organisation of the flow, similar periods and average amplitudes have been observed in the $(I,D)$ plane for unstable limit cycles in comparably small domains at smaller Reynolds numbers [Kreilos and Eckhardt, 2012]. Note however that the amplitude of the observed cycles was larger than what we display here by an order of magnitude. A very similar state has been observed at $R=400$ (not shown here).
@@
As the system size is increased to larger values $12\times 8$ (not shown) and $L_{x}\times L_{z}=18\times 12$, several changes occur in this saddle. The saddle becomes time invariant, both for kinetic energy (Fig. 7 (g)) and velocity and vorticity fields (Fig. 7 (e,f)): it is thus an actual saddle point. It still presents streamwise modulated streamwise velocity tubes that occupy the whole domain, however, these velocity tubes are less intense with $\max_{\text{space}}|u_{x}|=0.44\pm 0.01$. This leads to a steady value of kinetic energy of $E_{k}=0.2$, a third of the time average found in the MFU type system. Additionally, the streamwise modulation in the adjacent streamwise velocity tubes is of varicose type. The spatial organisation and average kinetic energy of this saddle is consistent with saddles computed at $R=400$ in domains of comparable size [Schneider et al., 2008]. The spatial variation of the streamwise vorticity is less sharp than in the MFU type systems and its amplitude is still moderate $|\omega_{x}|=\mathcal{O}(0.1)$ (Fig. 7 (f)). As the system size is further increased to $L_{x}\times L_{z}=36\times 24$, yet another change appears in the saddle which is reached after bisections: the state becomes localised in $z$ (Fig. 8 (e)) and modulated in $x$ to a smaller extent. the most intense streamwise velocity tube has $u_{x}\geq 0$ in our computation, with $\max_{\text{space}}|u_{x}|=0.78\pm 0.01$, it is flanked by pairs of negative then positive streamwise velocity tube of comparable, though decreasing intensity. The streamwise velocity then decays rapidly with $z$. Moreover the modulation of streamwise velocity is localised in $x$, leading to a streamwise localisation of streamwise vorticity (Fig. 8 (f)). The streamwise propagation of this modulation makes this saddle a travelling wave. This travelling wave is consistent with those observed at smaller Reynolds numbers $R=400$, in domains slightly larger in the spanwise direction and larger in the streamwise direction [Schneider et al., 2010].
@@
When the saddle state occupies the whole domain (at any given time for $L_{x}\times L_{z}=6\times 4$ and any time for $L_{x}\times L_{z}=12\times 8$ and $L_{x}\times L_{z}=18\times 12$), it has spatial symmetries: it is invariant under the operation $y\rightarrow-y$, $z\rightarrow-z$ $x\rightarrow-x+\delta x$ with $\delta x\lesssim 1$. The the domain of size $L_{x}\times L_{z}=6\times 4$, the time periodicity, the time periodicity can be viewed as a time symmetry. Moreover, advancing in time by half a period and applying the afore mentioned spatial transformation leads to a space time symmetry. Moreover, increasing the system size to $L_{x}\times L_{z}=12\times 8$ and to $L_{x}\times L_{z}=18\times 12$, means that the saddle has respectively two and three wavelengths of the same pattern in both $x$ and $z$ direction and thus respectively introduces the symmetry groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ and $\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$. The symmetry and the parity of velocity tubes numbers is broken in the larger domain $L_{x}\times L_{z}=36\times 24$, where the state is localised. This broken symmetry is accompanied by a breaking of the steadiness in $x$ of the saddle. Our trajectory led to a saddle which is a wave travelling in the $-\mathbf{e}_{x}$ direction. We note that this system can also sustain a travelling wave obtained under the operation $y\rightarrow-y$, $\mathbf{u}\rightarrow-\mathbf{u}$ [Schneider et al., 2010]. This conditions the direction of propagation as the celerity of the wave is related to $\int_{y,z}u_{x}\,{\rm d}y{\rm d}z\neq 0$ because of the spanwise localisation. More precisely, we find a carrier wave of wave number $n_{x}=4$ which is almost steady is space with an almost constant amplitude, with a travelling modulation at lower wavenumbers. We can observe this using equation (33, but this time at $y=-0.75$ and $z=12$) giving the rescaled phase of Fourier modes $n_{x}=4$ (part of the carrier) and $n_{x}=3$ (modulation) in figure 10 (b), this time at $z=12$ and $y=-0.75$ where $u_{z}$ is the most intense and most concentrated. The rescaled phase of mode $n_{x}=4$ is near constant and confirms us the spatial steadiness of the modulation. Meanwhile the monotonous decrease of the phase of mode $n_{x}=3$ confirms the travel of the pattern toward smaller $x$. The length $\Delta\psi=-24$ by which it has decreased in the time window is consistent with the observed movement of the pattern in the periodic domain. Note that this propagation occurs at constant spatially averaged kinetic energy (Fig. 8 (g)).
@@
We further our comparison of the respective properties of the reactive trajectories: the corresponding last state at the last stage on the one hand and the saddles which are the closest to the crossing point of the separatrix on the other hand. In particular, we perform this comparison in view of the properties of last states and saddles in many of observed noise induced transitions. This comparison highlights a striking difference. In most known stochastically forced systems, where the saddle is clearly identified, both types of states are very similar: in those cases, all the trajectories cross the separatrix in the neighbourhood of the saddle, in particular when the noise variance is smaller and trajectories concentrate in phase space. In the case of larger noise variance, trajectories tend to spread around the saddle, instead of all avoiding it. In our case, for all system sizes, the reactive trajectories cross the separatrix with velocity fields (Fig. 5 (b,c), Fig. 6 (b), Fig. 7 (c,d), Fig. 8 (c,d)) which remain fairly similar in scale and in velocity and vorticity amplitude. These amplitudes and scales are all clearly distinct from the nearest saddle on the separatrix (Fig. 5 (d,e), Fig. 6 (c,d), Fig. 7 (e,f), Fig. 8 (e,f)), whose stable manifold they cross. For all system sizes one common difference is that the amplitude of streamwise vortices is stronger ($\max_{\text{space}}|\omega_{x}|\lesssim 1$) on the trajectory as it crosses the separatrix as it is in the saddle. The spatial characteristic scale of said vortices is also smaller on the trajectory than it is on the saddle. The same can be said of the modulation of the streamwise velocity tubes. Other size dependent distinctions can be found. For the MFU type system, the trajectory cross the separatrix at a kinetic energy $E_{k}\simeq 0.02$ lower than that of the unstable limit cycle where $E_{k}\simeq 0.06$. This value increases very little with $\beta$ in the tested range. In the system of size $L_{x}\times L_{z}=12\times 8$ most reactive trajectories start localising in $z$ (not shown here). All trajectories localise in $z$ in the system of size $L_{x}\times L_{z}=18\times 12$, while the saddle is spread over the whole domain in those two systems. In the largest system of size $L_{x}\times L_{z}=36\times 24$, streamwise vorticity is more localised in $z$ than it is the case for the traveling wave. We will see in the next sections (§ 3.24.1) that this happens while all trajectories have the properties of concentrated trajectories for small enough forcing noise variance.
@@ 3.2 Statistical description of trajectories
We first present the average path and standard deviation around it in the plane formed by $\log(E_{k,y-z})$ and $\log(E_{k,x})$ (Fig. 11 (c)), in a system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$. The figure is generated using the same procedure as in [Rolland, 2022] § 3.1.2: we compute the average and standard deviation of $\log(E_{k,y-z})$, conditioned on the value of $\log(E_{k,x})$ on reactive trajectories. The tube formed by the average plus/minus the standard deviation thus indicates us where $66\%$ of paths go (using the reasonable assumption of a Gaussian spread around the mean). We compare the type of paths followed in the case of forcing on component $x$ (with $554$ paths) to paths obtained under divergence free forcing at increasing $\beta$ ($1600$ paths for $\beta=200000$, $687$ paths for $\beta=300000$, $738$ paths for $\beta=330000$, $718$ paths for $\beta=400000$, $740$ paths for $\beta=5000000$ and 755 paths for $\beta=600000$). These paths have an estimated probability of occurrence that range from $0.59\pm 0.01$ to $2.2\cdot 10^{-25}\pm 0.8\cdot 10^{-25}$ (which can be considered as rare). Each set of paths starts from a slightly different neighbourhood. Indeed, since $\mathcal{A}$ and $\mathcal{C}$ are defined using the average and standard deviation of kinetic energy conditioned on no transition occurring (§ 2.2.2), the smaller the energy injection rate is, the smaller are $E_{k-x}$ and $E_{k,y-z}$ at the beginning of the trajectories (Fig. 4). However, each group of paths (for decreased $\beta$ or changed forcing type) displays the same type of evolution, with clear convexity of the curve $\langle\log(E_{k,y-z})\rangle(\log(E_{k,x}))$. This confirms in a statistical manner the observation that the two types of forcing eventually lead to flows that follow the same type of paths as they build up toward turbulence, in that they cross the separatrix at similar locations, with similar flow patterns. Schematically, we first observe a growth of $\log(E_{k,x})$ then a growth of $\log(E_{k,y-z})$. This indicates that the velocity streaks would first receive energy from the forcing, mostly indirectly, and are in an externally sustained process. This part of the trajectory is similar to what is observed using deterministic targeted forcing [Rigas et al., 2021]. They do so until they are intense enough to naturally start the streaks instability and to go from an interplay of streaks and vortices sustained by energy injection to the self sustaining process of turbulence. In all cases, the trajectories are more and more concentrated around the corresponding average path as $\beta$ is increased and $\epsilon$ is decreased. For the lowest values of $\epsilon$, we sampled these trajectories from several independent AMS runs to ensure that this narrower variance was not an effect of a smaller variability on the first part of the path within an AMS run.
@@
We perform a similar observation, this time using the curves of all the reactive trajectories computed by means of AMS along with a relaxation from the fixed point computed using bisections started from a reactive trajectory in a system of size $L_{x}\times L_{z}=18\times 12$ at $\beta=80000$. The paths are represented in the three dimensional phase space formed by $\log(E_{k,x})$, $\log(E_{k,y-z})$ and the turbulent fraction $F$ (Fig. 11 (d)). This last quantity was added to discuss the effect of spatial localisation on the paths. With a probability of observing the transition at this energy injection rate at $\alpha=\mathcal{O}\left(10^{-5}\right)$, we observe a concentration of paths in phase space (this can even be observed at $\beta=6.25\cdot 10^{4}$ where $\alpha=\mathcal{O}\left(10^{-1}\right)$). In that case we first have the growth of $E_{k,x}$ (from $\log(E_{k,x})\lesssim-4.5$ to $\log(E_{k,x})\gtrsim-4$) followed by that of $E_{k,y-z}$ already observed in smaller systems, while $F=0$. This corresponds to the formation of a streamwise band where velocity streaks are intense enough, modulated and in interplay with spanwise localised streamwise vortices. Once this band is intense enough ($\log(E_{k,x}\gtrsim-4$, $\log(E_{k,y-z})\gtrsim-10$) the turbulent fraction $F$ become non-zero and turbulence expands in the $z$ direction. This is indicated by the correlated increase of $\log(E_{k,x})$, $\log(E_{k,y-z})$ and $F$.
@@
We note two peculiar facts. Firstly, while the paths are more and more concentrated around their average in the MFU sized system, these averages are distinct from one value of $\epsilon$ to another and the bands given by plus/minus the standard deviation do not overlap. We do not observe the typical narrower and narrower concentration around an instanton. We note however that these bands move in phase space as $\epsilon$ is decreased, which means that said instanton could be reached asymptotically. Secondly, we added to the graph of figure 11 (c) the path followed in the $\log(E_{k,y-z}),\log(E_{k,x})$ plane by the relaxation from the unstable limit cycle which is the nearest to the transition path to turbulence, as computed in section 3.1.3. At that size, the relaxation starts from values of $E_{k,x}$ close to what is found in turbulence and displays a burst of $E_{k,y-z}$ along with an overshoot of $E_{k,x}$: the path followed in that plane is distinct from reactive trajectories. This can be viewed as a corollary of the first observation. We similarly added the curve followed in the three dimensional space $\log(E_{k,x})$, $\log(E_{k,y-z})$, $F$ by the relaxation from the fixed point in the system of size $L_{x}\times L_{z}=18\times 12$ (Fig. 11 (d)). This path can be viewed as the relaxation path that would be followed by an instanton in this system. We note that none of the group of transitions paths follows that relaxation path. The relaxation starts from a low value of $E_{k,y-z}$ ($\log(E_{k,y-z})\simeq-9.7$) and moderate value of $E_{k,x}$ ($\log(E_{k,x})\simeq-3.9$). The zero value of $F$ is a consequence of the value of the threshold used to compute the turbulent fraction, as the saddle is delocalised in the whole domain. As the flow relaxes and follows the unstable direction, the departure to the edge state grows globally in the domain, not locally to the point at $\log(E_{k,x})\simeq-3.6$, $E_{k,y-z}=-8.3$, where values of $e$ (see Eq. (6)) are large enough everywhere to be above the detection threshold and the turbulent fraction jumps to $1$. The kinetic energy contained in both streaks and vortices then keep growing globally until the flow reaches transitional wall turbulence. Following both relaxations thus confirms what had been visualised on the field taken from examples.
@@
Figure 11: (a) Plot of a reactive trajectory obtained by means of AMS in the MFU type domain at $\beta=600000$ (blue line), relaxations toward laminar (red line) and turbulent flow (orange line) started from the reactive trajectory as well as the result of the bisections (green line) in the plane formed by the logarithm of kinetic energy on components $\log(E_{k_{,}x})$ and $\log(E_{k,y-z})$. An inset zooms on the area where the dynamics winds around the unstable limit cycle. (b) Plot of the same reactive trajectory, relaxation and result of bisections, in the plane $I,D$ where $I$ is the total energy injection and $D$ is the total dissipation (Eq. (9)), using the same colour code. The dashed line $D=I$ is added. An inset zooms on the area where the dynamics winds around the unstable limit cycle. (c) Average plus and minus standard deviation of the logarithm of the kinetic energy in $y-z$ components $\log(E_{k,y-z})$ conditioned on logarithm of the kinetic energy in $x$ component $\log(E_{k,x})$ obtained from reactive trajectories computed by means of AMS with a divergence free forcing and increasing $\beta$ as well as with forcing on the $x$ component. The path in the $(\log(E_{k,x}),\log(E_{k,y-z}))$ plane of a relaxation from the saddle toward turbulence is added in orange, with a dot indicating its starting point. (d) Three dimensional plot of 151 reactive trajectories obtained by means of AMS in the domain of size $L_{x}\times L_{z}=18\times 12$ at $\beta=80000$ (blue) as well as the relaxation from the fixed point (black) in the space formed by the logarithm of kinetic energy contained in components $\log(E_{k,x})$, $\log(E_{k,y-z})$ and turbulent fraction $F$.
@@
We then consider the distribution of durations of reactive trajectories, in a system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$. They are sampled using AMS with forcing on components and divergence free forcing. We display the distribution of durations rescaled by the sample average duration $\tau$ and the sample standard deviation $\sigma$ in figure 12 (a) in linear scale and in figure 12 (b) in logarithmic scale. In both cases, the distributions are compared to a normalised Gumbel distribution defined using the $\gamma$ Euler constant
$\mu(s)=\frac{\pi}{\sqrt{6}}\exp\left(-\left(\frac{\pi}{\sqrt{6}}s+\gamma+\exp\left(\frac{\pi}{\sqrt{6}}s+\gamma\right)\right)\right)\,,\,\gamma=0.577\ldots\,.$ (34)
@@
For one degree of freedom systems, in the limit of vanishing noise variance, it has been demonstrated that this is the distribution of duration [Cérou et al., 2013]. It has been observed that the sampled distribution is close to a Gumbel in more complex stochastic systems [Rolland et al., 2016, Rolland, 2018] when we have instanton phenomenology (with clear deviation when we do not, for instance when the saddle becomes degenerate [Rolland et al., 2016]). This distribution is also observed in systems not forced by noise [Rolland, 2022, Gomé et al., 2022]. In our case, we observe a very good match between the sampled distribution and a Gumbel, with small differences for $t\leq\tau$ (which was also observed for instanton phenomenology in extended gradient systems [Rolland et al., 2016]).

### v2

@@ 3.1.3 Bisections
Following the procedure presented in section 2.2.4, we perform relaxations and bisections using reactive trajectories computed by means of AMS. We will examine the type of unstable attractor (unstable fixed point, unstable limit cycle, unstable travelling wave) which is reached and describe its main features. The result of these bisections will be compared to reactive trajectories and the last state at the last step examined in the former subsection. We will check whether the system displays a standard instanton phenomenology, which would be characterised by greater and greater correspondence between $\mathbf{u}_{\rm last}$ and the result of bisections: this would mean the trajectories would cross the separatrix very near the saddle. The time series of these relaxations and the result of following bisections are added to the plots:
@@
This rescaling matches with a $x$ position. It is incremented by $L_{x}$ each time $\arg(\hat{u}_{z})$ goes from $2\pi$ to 00 and decremented by $-L_{x}$ each time $\arg(\hat{u}_{z})$ goes from 00 to $2\pi$. The succession of plateau separated by a rapid jump every period of the kinetic energy quantifies the fact that we periodically see the pattern in $\omega_{x}$ disappear and reappear at a distance $L_{x}/2$. This edge state in the form of an unstable limit cycle has not been specifically studied in the literature to our knowledge, possibly because of the smallness of the domain. However, a similar spatial organisation of the flow, similar periods and average amplitudes have been observed in the $(I,D)$ plane for unstable limit cycles in comparably small domains at smaller Reynolds numbers Kreilos and Eckhardt, (2012). Note however that the amplitude of the observed cycles was larger than what we display here by an order of magnitude. A very similar state has been observed at $R=400$ (not shown here).
@@
As the system size is increased to larger values $12\times 8$ (not shown) and $L_{x}\times L_{z}=18\times 12$, several changes occur in this saddle. The saddle becomes time invariant, both for kinetic energy (Fig. 7 (g)) and velocity and vorticity fields (Fig. 7 (e,f)): it is thus an actual saddle point. It still presents streamwise modulated streamwise velocity tubes that occupy the whole domain, however, these velocity tubes are less intense with $\max_{\text{space}}|u_{x}|=0.44\pm 0.01$. This leads to a steady value of kinetic energy of $E_{k}=0.2$, a third of the time average found in the MFU type system. Additionally, the streamwise modulation in the adjacent streamwise velocity tubes is of varicose type. The spatial organisation and average kinetic energy of this saddle is consistent with saddles computed at $R=400$ in domains of comparable size Schneider et al., (2008). The spatial variation of the streamwise vorticity is less sharp than in the MFU type systems and its amplitude is still moderate $|\omega_{x}|=\mathcal{O}(0.1)$ (Fig. 7 (f)). As the system size is further increased to $L_{x}\times L_{z}=36\times 24$, yet another change appears in the saddle which is reached after bisections: the state becomes localised in $z$ (Fig. 8 (e)) and modulated in $x$ to a smaller extent. the most intense streamwise velocity tube has $u_{x}\geq 0$ in our computation, with $\max_{\text{space}}|u_{x}|=0.78\pm 0.01$, it is flanked by pairs of negative then positive streamwise velocity tube of comparable, though decreasing intensity. The streamwise velocity then decays rapidly with $z$. Moreover the modulation of streamwise velocity is localised in $x$, leading to a streamwise localisation of streamwise vorticity (Fig. 8 (f)). The streamwise propagation of this modulation makes this saddle a travelling wave. This travelling wave is consistent with those observed at smaller Reynolds numbers $R=400$, in domains slightly larger in the spanwise direction and larger in the streamwise direction Schneider et al., (2010).
@@
When the saddle state occupies the whole domain (at any given time for $L_{x}\times L_{z}=6\times 4$ and any time for $L_{x}\times L_{z}=12\times 8$ and $L_{x}\times L_{z}=18\times 12$), it has spatial symmetries: it is invariant under the operation $y\rightarrow-y$, $z\rightarrow-z$ $x\rightarrow-x+\delta x$ with $\delta x\lesssim 1$. The the domain of size $L_{x}\times L_{z}=6\times 4$, the time periodicity, the time periodicity can be viewed as a time symmetry. Moreover, advancing in time by half a period and applying the afore mentioned spatial transformation leads to a space time symmetry. Moreover, increasing the system size to $L_{x}\times L_{z}=12\times 8$ and to $L_{x}\times L_{z}=18\times 12$, means that the saddle has respectively two and three wavelengths of the same pattern in both $x$ and $z$ direction and thus respectively introduces the symmetry groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ and $\mathbb{Z}/3\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}$. The symmetry and the parity of velocity tubes numbers is broken in the larger domain $L_{x}\times L_{z}=36\times 24$, where the state is localised. This broken symmetry is accompanied by a breaking of the steadiness in $x$ of the saddle. Our trajectory led to a saddle which is a wave travelling in the $-\mathbf{e}_{x}$ direction. We note that this system can also sustain a travelling wave obtained under the operation $y\rightarrow-y$, $\mathbf{u}\rightarrow-\mathbf{u}$ Schneider et al., (2010). This conditions the direction of propagation as the celerity of the wave is related to $\int_{y,z}u_{x}\,{\rm d}y{\rm d}z\neq 0$ because of the spanwise localisation. More precisely, we find a carrier wave of wave number $n_{x}=4$ which is almost steady is space with an almost constant amplitude, with a travelling modulation at lower wavenumbers. We can observe this using equation (16, but this time at $y=-0.75$ and $z=12$) giving the rescaled phase of Fourier modes $n_{x}=4$ (part of the carrier) and $n_{x}=3$ (modulation) in figure 10 (b), this time at $z=12$ and $y=-0.75$ where $u_{z}$ is the most intense and most concentrated. The rescaled phase of mode $n_{x}=4$ is near constant and confirms us the spatial steadiness of the modulation. Meanwhile the monotonous decrease of the phase of mode $n_{x}=3$ confirms the travel of the pattern toward smaller $x$. The length $\Delta\psi=-24$ by which it has decreased in the time window is consistent with the observed movement of the pattern in the periodic domain. Note that this propagation occurs at constant spatially averaged kinetic energy (Fig. 8 (g)).
@@
We further our comparison of the respective properties of the reactive trajectories: the corresponding last state at the last step on the one hand and the saddles which are the closest to the crossing point of the separatrix on the other hand. In particular, we perform this comparison in view of the properties of last states and saddles in many of observed noise induced transitions. This comparison highlights a striking difference. In most known stochastically forced systems, where the saddle is clearly identified, both types of states are very similar: in those cases, all the trajectories cross the separatrix in the neighbourhood of the saddle, in particular when the noise variance is smaller and trajectories concentrate in phase space. In the case of larger noise variance, trajectories tend to spread around the saddle, instead of all avoiding it. In our case, for all system sizes, the reactive trajectories cross the separatrix with velocity fields (Fig. 5 (b,c), Fig. 6 (b), Fig. 7 (c,d), Fig. 8 (c,d)) which remain fairly similar in scale and in velocity and vorticity amplitude. These amplitudes and scales are all clearly distinct from the nearest saddle on the separatrix (Fig. 5 (d,e), Fig. 6 (c,d), Fig. 7 (e,f), Fig. 8 (e,f)), whose stable manifold they cross. For all system sizes one common difference is that the amplitude of streamwise vortices is stronger ($\max_{\text{space}}|\omega_{x}|\lesssim 1$) on the trajectory as it crosses the separatrix as it is in the saddle. The spatial characteristic scale of said vortices is also smaller on the trajectory than it is on the saddle. The same can be said of the modulation of the streamwise velocity tubes. Other size dependent distinctions can be found. For the MFU type system, the trajectory cross the separatrix at a kinetic energy $E_{k}\simeq 0.02$ lower than that of the unstable limit cycle where $E_{k}\simeq 0.06$. This value increases very little with $\beta$ in the tested range. In the system of size $L_{x}\times L_{z}=12\times 8$ most reactive trajectories start localising in $z$ (not shown here). All trajectories localise in $z$ in the system of size $L_{x}\times L_{z}=18\times 12$, while the saddle is spread over the whole domain in those two systems. In the largest system of size $L_{x}\times L_{z}=36\times 24$, streamwise vorticity is more localised in $z$ than it is the case for the traveling wave. We will see in the next sections (§ 3.24.1) that this happens while all trajectories have the properties of concentrated trajectories for small enough forcing noise variance.
@@ 3.2 Statistical description of trajectories
We first present the average path and standard deviation around it in the plane formed by $\log(E_{k,y-z})$ and $\log(E_{k,x})$ (Fig. 12 (a)), in a system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$. The figure is generated using the same procedure as in Rolland, (2022) § 3.1.2: we compute the average and standard deviation of $\log(E_{k,y-z})$, conditioned on the value of $\log(E_{k,x})$ on reactive trajectories. The tube formed by the average plus/minus the standard deviation thus indicates us where $66\%$ of paths go (using the reasonable assumption of a Gaussian spread around the mean). We compare the type of paths followed in the case of forcing on component $x$ (with $554$ paths) to paths obtained under divergence free forcing at increasing $\beta$ ($1600$ paths for $\beta=200000$, $687$ paths for $\beta=300000$, $738$ paths for $\beta=330000$, $718$ paths for $\beta=400000$, $740$ paths for $\beta=5000000$ and 755 paths for $\beta=600000$). These paths have an estimated probability of occurrence that range from $0.59\pm 0.01$ to $2.2\cdot 10^{-25}\pm 0.8\cdot 10^{-25}$ (which can be considered as rare). Each set of paths starts from a slightly different neighbourhood. Indeed, since $\mathcal{A}$ and $\mathcal{C}$ are defined using the average and standard deviation of kinetic energy conditioned on no transition occurring (§ 2.2.2), the smaller the energy injection rate is, the smaller are $E_{k-x}$ and $E_{k,y-z}$ at the beginning of the trajectories (Fig. 4). However, each group of paths (for decreased $\beta$ or changed forcing type) displays the same type of evolution, with clear convexity of the curve $\langle\log(E_{k,y-z})\rangle(\log(E_{k,x}))$. This confirms in a statistical manner the observation that the two types of forcing eventually lead to flows that follow the same type of paths as they build up toward turbulence, in that they cross the separatrix at similar locations, with similar flow patterns. Schematically, we first observe a growth of $\log(E_{k,x})$ then a growth of $\log(E_{k,y-z})$. This indicates that the velocity streaks would first receive energy from the forcing, mostly indirectly, and are in an externally sustained process. This part of the trajectory is similar to what is observed using deterministic targeted forcing Rigas et al., (2021). They do so until they are intense enough to naturally start the streaks instability and to go from an interplay of streaks and vortices sustained by energy injection to the self sustaining process of turbulence. In all cases, the trajectories are more and more concentrated around the corresponding average path as $\beta$ is increased and $\epsilon$ is decreased. For the lowest values of $\epsilon$, we sampled these trajectories from several independent AMS runs to ensure that this narrower variance was not an effect of a smaller variability on the first part of the path within an AMS run.
@@
We perform a similar observation, this time using the curves of all the reactive trajectories computed by means of AMS along with a relaxation from the fixed point computed using bisections started from a reactive trajectory in a system of size $L_{x}\times L_{z}=18\times 12$ at $\beta=80000$. The paths are represented in the three dimensional phase space formed by $\log(E_{k,x})$, $\log(E_{k,y-z})$ and the turbulent fraction $F$ (Fig. 12 (b)). This last quantity was added to discuss the effect of spatial localisation on the paths. With a probability of observing the transition at this energy injection rate at $\alpha=\mathcal{O}\left(10^{-5}\right)$, we observe a concentration of paths in phase space (this can even be observed at $\beta=6.25\cdot 10^{4}$ where $\alpha=\mathcal{O}\left(10^{-1}\right)$). In that case we first have the growth of $E_{k,x}$ (from $\log(E_{k,x})\lesssim-4.5$ to $\log(E_{k,x})\gtrsim-4$) followed by that of $E_{k,y-z}$ already observed in smaller systems, while $F=0$. This corresponds to the formation of a streamwise band where velocity streaks are intense enough, modulated and in interplay with spanwise localised streamwise vortices. Once this band is intense enough ($\log(E_{k,x}\gtrsim-4$, $\log(E_{k,y-z})\gtrsim-10$) the turbulent fraction $F$ become non-zero and turbulence expands in the $z$ direction. This is indicated by the correlated increase of $\log(E_{k,x})$, $\log(E_{k,y-z})$ and $F$.
@@
We note two peculiar facts. Firstly, while the paths are more and more concentrated around their average in the MFU sized system, these averages are distinct from one value of $\epsilon$ to another and the bands given by plus/minus the standard deviation do not overlap. We do not observe the typical narrower and narrower concentration around an instanton. We note however that these bands move in phase space as $\epsilon$ is decreased, which means that said instanton could be reached asymptotically. Secondly, we added to the graph of figure 12 (a) the path followed in the $\log(E_{k,y-z}),\log(E_{k,x})$ plane by the relaxation from the unstable limit cycle which is the nearest to the transition path to turbulence, as computed in section 3.1.3. At that size, the relaxation starts from values of $E_{k,x}$ close to what is found in turbulence and displays a burst of $E_{k,y-z}$ along with an overshoot of $E_{k,x}$: the path followed in that plane is distinct from reactive trajectories.
This can be viewed as a corollary of the first observation. We similarly added the curve followed in the three dimensional space $\log(E_{k,x})$, $\log(E_{k,y-z})$, $F$ by the relaxation from the fixed point in the system of size $L_{x}\times L_{z}=18\times 12$ (Fig. 12 (b)). This path can be viewed as the relaxation path that would be followed by an instanton in this system. We note that none of the group of transitions paths follows that relaxation path. The relaxation starts from a low value of $E_{k,y-z}$ ($\log(E_{k,y-z})\simeq-9.7$) and moderate value of $E_{k,x}$ ($\log(E_{k,x})\simeq-3.9$). The zero value of $F$ is a consequence of the value of the threshold used to compute the turbulent fraction, as the saddle is delocalised in the whole domain. As the flow relaxes and follows the unstable direction, the departure to the edge state grows globally in the domain, not locally to the point at $\log(E_{k,x})\simeq-3.6$, $E_{k,y-z}=-8.3$, where values of $e$ (see Eq. (6)) are large enough everywhere to be above the detection threshold and the turbulent fraction jumps to $1$. The kinetic energy contained in both streaks and vortices then keep growing globally until the flow reaches transitional wall turbulence. Following both relaxations thus confirms what had been visualised on the field taken from examples.
@@
Figure 11: (a) Plot of a reactive trajectory obtained by means of AMS in the MFU type domain at $\beta=600000$ (blue line), relaxations toward laminar (red line) and turbulent flow (orange line) started from the reactive trajectory as well as the result of the bisections (green line) in the plane formed by the logarithm of kinetic energy on components $\log(E_{k_{,}x})$ and $\log(E_{k,y-z})$. An inset zooms on the area where the dynamics winds around the unstable limit cycle. (b) Plot of the same reactive trajectory, relaxation and result of bisections, in the plane $I,D$ where $I$ is the total energy injection and $D$ is the total dissipation (Eq. (10)), using the same colour code. The dashed line $D=I$ is added. An inset zooms on the area where the dynamics winds around the unstable limit cycle.
Refer to caption (b) Refer to caption
Figure 12: (a) Average plus and minus standard deviation of the logarithm of the kinetic energy in $y-z$ components $\log(E_{k,y-z})$ conditioned on logarithm of the kinetic energy in $x$ component $\log(E_{k,x})$ obtained from reactive trajectories computed by means of AMS with a divergence free forcing and increasing $\beta$ as well as with forcing on the $x$ component (termed wcx in the caption). The path in the $(\log(E_{k,x}),\log(E_{k,y-z}))$ plane of a relaxation from the saddle toward turbulence is added in orange, with a dot indicating its starting point. (b) Three dimensional plot of 151 reactive trajectories obtained by means of AMS in the domain of size $L_{x}\times L_{z}=18\times 12$ at $\beta=80000$ (blue) as well as the relaxation from the fixed point (black) in the space formed by the logarithm of kinetic energy contained in components $\log(E_{k,x})$, $\log(E_{k,y-z})$ and turbulent fraction $F$.
@@
We then consider the distribution of durations of reactive trajectories, in a system of size $L_{x}\times L_{z}=6\times 4$ at $R=500$. They are sampled using AMS with forcing on components and divergence free forcing. We display the distribution of durations rescaled by the sample average duration $\tau$ and the sample standard deviation $\sigma$ in figure 13 (a) in linear scale and in figure 13 (b) in logarithmic scale. In both cases, the distributions are compared to a normalised Gumbel distribution defined using the Euler constant $\gamma=0.577\ldots$
$\mu(s)=\frac{\pi}{\sqrt{6}}\exp\left(-\left(\frac{\pi}{\sqrt{6}}s+\gamma+\exp\left(\frac{\pi}{\sqrt{6}}s+\gamma\right)\right)\right)\,.$ (17)
@@
For one degree of freedom systems, in the limit of vanishing noise variance, it has been demonstrated that this is the distribution of duration Cérou et al., (2013). It has been observed that the sampled distribution is close to a Gumbel in more complex stochastic systems Rolland et al., (2016); Rolland, (2018) when we have instanton phenomenology (with clear deviation when we do not, for instance when the saddle becomes degenerate Rolland et al., (2016)). This distribution is also observed in systems not forced by noise Rolland, (2022); Gomé et al., (2022). In our case, we observe a very good match between the sampled distribution and a Gumbel, with small differences for $t\leq\tau$ (which was also observed for instanton phenomenology in extended gradient systems Rolland et al., (2016)).